# Towards Multi-agent Reinforcement Learning based Traffic Signal Control through Spatio-temporal Hypergraphs

### v1

@@ Abstract
Traffic signal control systems (TSCSs) are integral to intelligent traffic management, fostering efficient vehicle flow. Traditional approaches often simplify road networks into standard graphs, which results in a failure to consider the dynamic nature of traffic data at neighboring intersections, thereby neglecting higher-order interconnections necessary for real-time control. To address this, we propose a novel TSCS framework to realize intelligent traffic control. This framework collaborates with multiple neighboring edge computing servers to collect traffic information across the road network. To elevate the efficiency of traffic signal control, we have crafted a multi-agent soft actor-critic (MA-SAC) reinforcement learning algorithm. Within this algorithm, individual agents are deployed at each intersection with a mandate to optimize traffic flow across the entire road network collectively. Furthermore, we introduce hypergraph learning into the critic network of MA-SAC to enable the spatio-temporal interactions from multiple intersections in the road network. This method fuses hypergraph and spatio-temporal graph structures to encode traffic data and capture the complex spatial and temporal correlations between multiple intersections. Our empirical evaluation, tested on varied datasets, demonstrates the superiority of our framework in minimizing average vehicle travel times and sustaining high-throughput performance. This work facilitates the development of more intelligent and reactive urban traffic management solutions.
@@
Traffic signal control, hypergraph learning, deep reinforcement learning
@@ I Introduction
Intelligent TSCS is a viable solution to dynamically control the traffic signal cost-effectively [3, 4]. Empowered by the state-of-the-art technologies including information sensing, data analysis, and optimization algorithms, the intelligent system is expected to automatically adjust the timing of traffic signal control according to real-time traffic conditions. Intelligent TSCSs offer significant advantages over static or manually controlled methods. These systems leverage real-time and historical data collected by monitoring sensors and cameras to make superior decision-making. Additionally, they have the ability to optimize the timing and cycle of traffic signal based on the current traffic flow, thereby improving the road condition.
@@
Reinforcement learning algorithms have been widely studied to achieve effective traffic signal control [8]. Specifically, the problem of controlling traffic signals at multiple intersections can be formulated as a Markov decision process (MDP), wherein training can yield a policy for selecting the optimal action at each state [9]. Reinforcement learning algorithms can be applied to solve the MDP by learning the optimal decision on traffic signal control through value functions and policy optimization methods. Among these algorithms, soft actor-critic (SAC) is a potential candidate to extract the traffic information due to the inherent stochasticity of the policy with entropy regularization. State-of-the-art work has demonstrated the effectiveness of the SAC-based method in training optimally control decisions for traffic signal. The primary achievement includes attaining a consistently high traffic throughput while substantially reducing the average travel time for all vehicles [10, 11].
@@ III-B Problem Description
  • Reward $\mathcal{R}$: Agent $i$ obtains a reward $r_{i}^{t}$ by reward function $\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{N}\rightarrow\mathcal{R}$. In this paper, the agent’s objective is to minimize the travel time for all vehicles in the road network.
@@
For the given road traffic network, each traffic signal agent $i$ at time $t$ makes action $a_{i}^{t}$ with reward $r_{i}^{t}$. The objective is to approximate the total reward $\sum_{\forall i}r_{i}^{t}$ by minimizing the loss. It is worth noting that, besides the loss defined in reinforcement learning algorithm, we additionally include the hypergraph reconstruction loss (See $\textit{loss}_{i}^{\textit{RECON}}$ in Figure 2) to enhance the training performance of the proposed framework. More details will be described in Section IV.
@@ IV-A Framework
Figure 2 illustrates an overview of our proposed framework. In this framework, the road information at each intersection is collected from the MEC server. The role of the agent is to make an efficient control decision on the traffic signal phase for the respective intersection. Edge intelligence is used to enhance the data processing performance by sharing the necessary information among various MEC servers. The construction of hypergraph involves incorporating multiple surrounding agents to acquire the spatial and temporal structure information, i.e., $o_{t}$ includes the state information of all intersections at time $t$. Meanwhile, the training process considers the historical state information as incorporating temporal dependency, i.e., ($o_{t}$, $o_{t-\Delta t}$) that covers the information at time $t$ and $t-\Delta t$.
@@ IV-C Spatio-temporal Hypergraph Construction
Owing to the dynamic nature of traffic flow and road conditions, employing initial hypergraph structures alone would overlook the dynamic modifications of such structures from adjusted feature embedding. As a result, some important implicit relationships may not be directly reflected in the inherent structural framework of hypergraphs. Therefore, it is necessary to dynamically modify the hypergraph structure during the process of model optimization to adapt to the changes of traffic conditions in the road network. We propose a dynamic learning process that generates two types of hyperedges: spatial hyperedges that capture heterogeneity between traffic light agents, encoding relations among different intersections in one timestamp, and temporal hyperedges that scrutinize interactivity of historical traffic flow, modeling the continuous interaction of road intersections in consecutive timestamps. During the dynamic hyperedges generation process, we introduce two kinds of nodes: master node and slave node. A master node $\dot{v}\in\mathcal{V}$ acts as an anchor when generating the hyperedge $e\left(\dot{v}\right)\in\mathcal{E}$, combining with a set of slave nodes $S\left(\dot{v}\right)=\left\{\hat{v}\right\}$ to collectively make up the hyperedge.
For each master node $\dot{v}^{t}_{i}\in\emph{V}^{t}$ in spatio-temporal hypergraph $\mathcal{G}^{t}$ at timestamp $t$, the spatial hyperedge $e_{spa}\left(\dot{v}_{i}^{t}\right)\in\mathcal{E}^{t}$ can be generated based on the reconstruction of the master node $\dot{v}^{t}_{i}$ and the spatial candidate slave node set $\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)=\left\{v|v\in\emph{V}^{t},v\neq\dot{v}^{t}_{i}\right\}$, which is denoted as:
@@
where ${c}_{spa}\left(\dot{v}^{t}_{i}\right)$ denotes the spatial reconstruction error. $\left\|\cdot\right\|_{2}$ denotes the $l2$ norm of the vector. $\emph{H}_{\mathcal{G}}^{t}\left(\dot{v}^{t}_{i}\right)$ and $\emph{H}_{\mathcal{G}}^{t}\left(\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)\right)$ are node feature matrices of the master node and the spatial candidate slave node set respectively. $\theta_{spa}$ is a specific trainable projection matrix when generating the spatial hyperedge $e_{spa}\left(\dot{v}_{i}^{t}\right)\in\mathcal{E}^{t}$. $\emph{p}^{spa}_{\dot{v}^{t}_{i}}\in\mathcal{R}^{\left(N-1\right)}$ denotes the trainable reconstruction coefficient vector, with each element $\emph{p}^{spa}_{\dot{v}^{t}_{i}}\left(v\right)$ representing the learned reconstruction coefficient of each node $v\in\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)$ relative to the master node $\dot{v}^{t}_{i}$. According to $\emph{p}^{spa}_{\dot{v}^{t}_{i}}$, the nodes in the spatial candidate slave node set $\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)$ with reconstruction coefficient larger than the threshold $\zeta$ are selected to generate a spatial hyperedge of the master node $\dot{v}^{t}_{i}$ (See the connected through green solid lines in Figure 2). In contrast, unselected nodes with reconstructions coefficient values lower than $\zeta$ are connected by dotted lines. This facilitates the dynamic learning of the reconstruction coefficients between the master node and each other node within the road network. By comparing with threshold $\zeta$, we are able to dynamically filter nodes more appropriate for interactions, not merely acquiring the correlation coefficients of neighbor nodes to the master node. Consequently, our approach offers greater adaptability than traditional graph-based methods when it comes to handling the interplay of spatio-temporal information between intersections. This allows for a refined and contextual response to the ever-changing dynamics of traffic networks. $S^{spa}\left(\dot{v}^{t}_{i}\right)$ is denoted as:
@@
Similarly, as shown in Fig. 2, temporal slave nodes are encircled from $\emph{V}^{t-1}$ to form the temporal hyperedge $e_{tem}\left(\dot{v}_{i}^{t}\right)$ with the master node based on the trainable reconstruction coefficient vector $\emph{p}^{tem}_{\dot{v}^{t}_{i}}\in\mathcal{R}^{N}$. Overall, the loss of hyperedges generation in one timestamp is defined as:
@@ IV-D1 Hyperedge Embedding Updating
Let Re represents the correlation between all nodes (within and not within the slave node set $S^{spa}(\dot{v}^{t}_{i})$) in the road network and the master node $\dot{v}^{t}_{i}$ at timestamps t, with entries are as below:
@@ V-A Experimental Settings
We introduce four synthetic traffic datasets ($\textit{Unidirect}_{\textit{6$\times$6}}$, $\textit{Bidirect}_{\textit{6$\times$6}}$, $\textit{Unidirect}_{\textit{10$\times$10}}$, and $\textit{Bidirect}_{\textit{10$\times$10}}$ ) and two real-world traffic datasets ($D_{\textit{Hangzhou}}$, and $D_{\textit{Jinan}}$) to validate the robustness of our proposed method. The details of them are as below:
@@ V-B Baseline Methods
We verify the performance of our proposed hypergraph-based deep reinforcement learning method, referred to as HG-DRL, with 6 comparative methods. These methods include the traditional transportation methods, reinforcement learning methods, and some that incorporate graph learning. The details of the baseline methods are described as below:
@@
  • PressLight [24]: Utilizing the concept of MaxPressure and deep reinforcement learning to effectively optimize the pressure at each intersection. This method is designed to solve the multi-intersection signal control problems.
  • MPLight [19]: This method utilizes the concept of pressure to achieve signal coordination at the regional level in a reinforcement learning-based approach, while also employing a network structure specifically designed to handle unbalanced traffic flow.
  • CoLight [28]: A graph attention network is introduced in the reinforcement learning setting of multi-intersection traffic signal control to enhance coordination and decision-making.
  • GCN-SAC: We design this method for conducting ablation experiments to validate the performance between standard graphs and hypergraphs. In the reinforcement learning module, we set GCN-SAC to be the identical as HG-DRL, while in GCN-SAC, graph attention network (GAT) is used to facilitate information interaction between adjacent nodes.
@@ V-D Performance Evaluation
Figure 4: Impact of hypergraph loss coefficient $\beta$.
@@
V-E Sensitive Analysis

### v2

@@ Abstract
Traffic signal control systems (TSCSs) are integral to intelligent traffic management, fostering efficient vehicle flow. Traditional approaches often simplify road networks into standard graphs, which results in a failure to consider the dynamic nature of traffic data at neighboring intersections, thereby neglecting higher-order interconnections necessary for real-time control. To address this, we propose a novel TSCS framework to realize intelligent traffic control. This framework collaborates with multiple neighboring edge computing servers to collect traffic information across the road network. To elevate the efficiency of traffic signal control, we have crafted a multi-agent soft actor-critic (MA-SAC) reinforcement learning algorithm. Within this algorithm, individual agents are deployed at each intersection with a mandate to optimize traffic flow across the road network collectively. Furthermore, we introduce hypergraph learning into the critic network of MA-SAC to enable the spatio-temporal interactions from multiple intersections in the road network. This method fuses hypergraph and spatio-temporal graph structures to encode traffic data and capture the complex spatio-temporal correlations between multiple intersections. Our empirical evaluation, tested on varied datasets, demonstrates the superiority of our framework in minimizing average vehicle travel times and sustaining high-throughput performance. This work facilitates the development of more intelligent urban traffic management solutions. We release the code to support the reproducibility of this work at https://github.com/Edun-Eyes/TSC.
@@
Traffic signal control, mobile data mining, multi-access edge computing, hypergraph learning, deep reinforcement learning
@@ I Introduction
Intelligent TSCS is a viable solution to dynamically control the traffic signal cost-effectively [3, 4, 5]. Empowered by the state-of-the-art technologies including information sensing, data analysis, and optimization algorithms, the intelligent system is expected to automatically adjust the timing of traffic signal control according to real-time traffic conditions. Intelligent TSCSs offer significant advantages over static or manually controlled methods. These systems leverage real-time and historical data collected by monitoring sensors and cameras to make superior decision-making. Additionally, they have the ability to optimize the timing and cycle of traffic signal based on the current traffic flow, thereby improving the road condition.
@@
Reinforcement learning algorithms have been widely studied to achieve effective traffic signal control [9, 10, 11]. Specifically, the problem of controlling traffic signals at multiple intersections can be formulated as a Markov decision process (MDP), wherein training can yield a policy for selecting the optimal action at each state [12]. Reinforcement learning algorithms can be applied to solve the MDP by learning the optimal decision on traffic signal control through value functions and policy optimization methods. Among these algorithms, soft actor-critic (SAC) is a potential candidate to extract the traffic information due to the inherent stochasticity of the policy with entropy regularization. State-of-the-art work has demonstrated the effectiveness of the SAC-based method in training optimally control decisions for traffic signal. The primary achievement includes attaining a consistently high traffic throughput while substantially reducing the average travel time for all vehicles [13, 14].
@@ II-C Hypergraph Learning
II-D Discussion
Recent researches like CoLight [31] have explored the application of reinforcement learning in traffic signal control by incorporating standard pairwise graph learning to model road network structures to enhance signal control across multiple intersections. However, these approaches primarily focus on the traffic data from first-order neighbors at each intersection, which neglects the impact of more distant intersections and failing to consider the multi-edge relationships among various nodes. Furthermore, real-world datasets exhibit dynamic fluctuations in traffic flow that create higher-order correlations between intersections, which is critical for standard graph learning models to adequately capture. This limitation poses challenges in effectively managing complex road networks with diverse topological structures.
@@ III-B Problem Description
  • Reward $\mathcal{R}$: Agent $i$ obtains a reward $r_{i}^{t}$ by reward function $\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{N}\rightarrow\mathcal{R}$. In this paper, the agent’s objective is to minimize the travel time for all vehicles in the road network. We define the reward for each intersection $n$ as:
    $r_{n}^{t}=\text{-}\sum_{l}{u_{n,l}^{t}}$ (1)
    where $u_{n,l}^{t}$ is the queue length on the approaching lane $l$ at time $t$. For each intersection, the waiting time for vehicles is influenced by the length of the queue in the lane. As the queue lengthens, the waiting time for vehicles increases, which results in a rise in overall travel time.
@@
For the given road traffic network, each traffic signal agent $i$ at time $t$ makes action $a_{i}^{t}$ with reward $r_{i}^{t}$. The objective is to approximate the total reward $\sum_{\forall i}r_{i}^{t}$ by minimizing the loss. It is worth noting that, besides the loss defined in reinforcement learning algorithm, we additionally include the hypergraph reconstruction loss (See $\textit{loss}_{i}^{\textit{RECON}}$ in Figure 2) to enhance the training performance of the proposed framework. We assume all agents are trained jointly while parameter sharing is not implemented due to the significant variation in the influence of each intersection’s neighbors on the target intersection [31]. More details will be described in Section IV.
@@ IV-A Framework
Figure 2 illustrates an overview of our proposed framework. In this framework, the road information at each intersection is collected from the MEC server. The role of the agent is to make an efficient control decision on the traffic signal phase for the respective intersection. Edge intelligence is used to enhance the data processing performance by sharing the necessary information among various MEC servers. Edge intelligence enables computation close to the data source, thereby significantly reducing data transmission latency and bandwidth requirements. Moreover, edge intelligence allows each agent to rapidly process traffic data locally, facilitating real-time traffic information sharing between intersections and the training of multiple intelligent agents deployed at intersections. This collaborative mechanism can better address the complexities of traffic flow, leading to more efficient traffic signal control. .
The construction of hypergraph involves incorporating multiple surrounding agents to acquire the spatial and temporal structure information, i.e., $o_{t}$ includes the state information of all intersections at time $t$. Meanwhile, the training process considers the historical state information as incorporating temporal dependency, i.e., ($o_{t}$, $o_{t-\Delta t}$) that covers the information at time $t$ and $t-\Delta t$.
@@ IV-C Spatio-temporal Hypergraph Construction
Owing to the dynamic nature of traffic flow and road conditions, employing initial hypergraph structures alone would overlook the dynamic modifications of such structures from adjusted feature embedding. As a result, some important implicit relationships may not be directly reflected in the inherent structural framework of hypergraphs. Therefore, it is necessary to dynamically modify the hypergraph structure during the process of model optimization to adapt to the changes of traffic conditions in the road network. We propose a dynamic learning process that generates two types of hyperedges: spatial hyperedges that capture heterogeneity between traffic light agents, encoding relations among different intersections in one timestamp, and temporal hyperedges that scrutinize interactivity of historical traffic flow, modeling the continuous interaction of road intersections in consecutive timestamps. During the dynamic hyperedges generation process, we introduce two kinds of nodes: master node and candidate node. A master node $\dot{v}\in\mathcal{V}$ acts as an anchor when generating the hyperedge $e\left(\dot{v}\right)\in\mathcal{E}$, combining with a set of candidate nodes $S\left(\dot{v}\right)=\left\{\hat{v}\right\}$ to collectively make up the hyperedge.
For each master node $\dot{v}^{t}_{i}\in\emph{V}^{t}$ in spatio-temporal hypergraph $\mathcal{G}^{t}$ at timestamp $t$, the spatial hyperedge $e_{spa}\left(\dot{v}_{i}^{t}\right)\in\mathcal{E}^{t}$ can be generated based on the reconstruction of the master node $\dot{v}^{t}_{i}$ and the spatial candidate node set $\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)=\left\{v|v\in\emph{V}^{t},v\neq\dot{v}^{t}_{i}\right\}$, which is denoted as:
@@
where ${c}_{spa}\left(\dot{v}^{t}_{i}\right)$ denotes the spatial reconstruction error. $\left\|\cdot\right\|_{2}$ denotes the $l2$ norm of the vector. $\emph{H}_{\mathcal{G}}^{t}\left(\dot{v}^{t}_{i}\right)$ and $\emph{H}_{\mathcal{G}}^{t}\left(\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)\right)$ are node feature matrices of the master node and the spatial candidate node set respectively. $\theta_{spa}$ is a specific trainable projection matrix when generating the spatial hyperedge $e_{spa}\left(\dot{v}_{i}^{t}\right)\in\mathcal{E}^{t}$. $\emph{p}^{spa}_{\dot{v}^{t}_{i}}\in\mathcal{R}^{\left(N-1\right)}$ denotes the trainable reconstruction coefficient vector, with each element $\emph{p}^{spa}_{\dot{v}^{t}_{i}}\left(v\right)$ representing the learned reconstruction coefficient of each node $v\in\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)$ relative to the master node $\dot{v}^{t}_{i}$. According to $\emph{p}^{spa}_{\dot{v}^{t}_{i}}$, the nodes in the spatial candidate node set $\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)$ with reconstruction coefficient larger than the threshold $\zeta$ are selected to generate a spatial hyperedge of the master node $\dot{v}^{t}_{i}$ (See the connected through green solid lines in Figure 2). In contrast, unselected nodes with reconstructions coefficient values lower than $\zeta$ are connected by dotted lines. This facilitates the dynamic learning of the reconstruction coefficients between the master node and each other node within the road network. By comparing with threshold $\zeta$, we are able to dynamically filter nodes more appropriate for interactions, not merely acquiring the correlation coefficients of neighbor nodes to the master node. Consequently, our approach offers greater adaptability than traditional graph-based methods when it comes to handling the interplay of spatio-temporal information between intersections. This allows for a refined and contextual response to the ever-changing dynamics of traffic networks. $S^{spa}\left(\dot{v}^{t}_{i}\right)$ is denoted as:
@@
Similarly, as shown in Fig. 2, temporal candidate nodes are encircled from $\emph{V}^{t-1}$ to form the temporal hyperedge $e_{tem}\left(\dot{v}_{i}^{t}\right)$ with the master node based on the trainable reconstruction coefficient vector $\emph{p}^{tem}_{\dot{v}^{t}_{i}}\in\mathcal{R}^{N}$. Overall, the loss of hyperedges generation in one timestamp is defined as:
@@ IV-D1 Hyperedge Embedding Updating
Let Re represents the correlation between all nodes (within and not within the candidate node set $S^{spa}(\dot{v}^{t}_{i})$) in the road network and the master node $\dot{v}^{t}_{i}$ at timestamps t, with entries are as below:
@@ V-A Experimental Settings
We introduce four synthetic traffic datasets ($\textit{Unidirect}_{\textit{6$\times$6}}$, $\textit{Bidirect}_{\textit{6$\times$6}}$, $\textit{Unidirect}_{\textit{10$\times$10}}$, and $\textit{Bidirect}_{\textit{10$\times$10}}$ ) and two real-world traffic datasets ($D_{\textit{Hangzhou}}$ and $D_{\textit{Jinan}}$) to validate the robustness of our proposed method. The details of them are as below:
@@ V-B Baseline Methods
We verify the performance of our proposed hypergraph-based deep reinforcement learning method, referred to as HG-DRL, with 6 comparative methods. These methods include the traditional transportation methods, reinforcement learning methods, and some that incorporate graph learning. To ensure a fair comparison, we utilize the optimal parameters configurations for various datasets as specified in the original open sourcecodes of the baseline methods. The details of the baseline methods are described as below:
@@
  • PressLight [27]: Utilizing the concept of MaxPressure and deep reinforcement learning to effectively optimize the pressure at each intersection. This method is designed to solve the multi-intersection signal control problems. The batch size is 20, episodes is 100, learning rate is 0,001, replay memory is 3000, discount rate is 0.8, and Adam is used as optimizer.
  • MPLight [22]: This method utilizes the concept of pressure to achieve signal coordination at the regional level in a reinforcement learning-based approach, while also employing a network structure specifically designed to handle unbalanced traffic flow. The key parameters are configured to match those used in PressLight.
  • CoLight [31]: A graph attention network is introduced in the reinforcement learning setting of multi-intersection traffic signal control to enhance coordination and decision-making. The replay memory is 1000, while RMSProp is used as optimizer. Other main parameters are set as the same as those in Presslight.
  • GCN-SAC: We design this method for conducting ablation experiments to validate the performance between standard graphs and hypergraphs. In the reinforcement learning module, we set GCN-SAC to be the identical as HG-DRL, while in GCN-SAC, graph attention network (GAT) is used to facilitate information interaction between adjacent nodes. The key parameters are configured to match those used in HG-DRL.
@@ V-D Performance Evaluation
In terms of the learning curves for different methods, our proposed method HG-DRL demonstrates enhanced post-convergence stability compared to other baseline methods. This stability arises from HG-DRL’s ability to dynamically construct hyperedges, enabling efficient capture and representation of complex spatio-temporal correlations inherent in traffic signal control systems. For example, CoLight demonstrates notable fluctuations in result values across a broad range of episodes, whereas our proposed method maintains consistent stability within the same range. This further substantiates the capability of HG-DRL to sustain stability across a wide range while achieving optimal performance. These findings align with the trends observed in the throughput evolution, as shown in Figure 4.
(a) $D_{\textit{Hangzhou}}$
(b) $D_{\textit{Jinan}}$
Figure 3: Learning curves for different methods.
(a) $D_{\textit{Hangzhou}}$
(b) $D_{\textit{Jinan}}$
Figure 4: Throughput evolution for different methods.
@@
Figure 6: Impact of hypergraph loss coefficient $\beta$
@@
V-E Ablation Study
We conduct the ablation study to validate the importance of hypergraph and DRL across three different datasets as shown in Table IV. In this table, “w/o HG” indicates the replacement of hypergraph with traditional pairwise GAT, while “w/o DRL” removes the DRL and installs regular actor-critic method. The results demonstrate that the hypergraph module enhances the average travel time by at most 16% compared with the one without it, which verifies that the dynamic construction of spatial and temporal hyperedges effectively captures the spatio-temporal dependencies among traffic signals within road network. Meanwhile, HG-DRL with DRL can outperform that with actor-critic in terms of travel time up to 10%. This advantage primarily attributes to the ability of MA-SAC to adaptively learn optimal policies by analyzing the interaction between traffic signal timing and road traffic flow dynamics.
TABLE IV: Results of ablation study in terms of average travel time
  Method $D_{\textit{Hangzhou}}$ $D_{\textit{Jinan}}$ $\textit{Bidirect}_{\textit{6$\times$6}}$
  w/o HG 336.15 332.36 172.31
w/o DRL 305.4 310.60 175.4
HG-DRL (Ours) 295.71 278.22 167.77
V-F Sensitive Analysis